A storage tank acquired at the beginning of the fiscal year at a cost of $72,000 has an estimated residual value of $4,400 and an estimated useful life of 4 years.

a. Determine the amount of annual depreciation by the straight-line method.
$

b. Determine the amount of depreciation for the first and second years computed by the double-declining-balance method. Do not round the double-declining balance rate. If required, round your answers to the nearest dollar.

Depreciation

Year 1$

Year 2$
Answer #1
a
Cost 72000
Less: Residual value 4400
Depreciable cost 67600
Divide by Estimated useful life 4
Annual depreciation 16900
b
Double-declining balance rate 50% =1/4*2
Depreciation:
Year 1 36000 =72000*50%
Year 2 18000 =(72000-36000)*50%